OpenStudy (anonymous):

Please, please, please help me out with this!!! Find the sum of the first six terms of the geometric sequence for which a2= 0.7 and a3= 0.49.

OpenStudy (anonymous):

well the series is geometric based on the information a general term is \[a_{n} = a_{1} r^{n-1}\] a1 = 1st term, r = common ratio so start by finding the common ratio 2nd term \[0.7 = ar\] 3rd term \[0.49 = ar^2... or..... 0.49 = (ar) \times r ...or...... 0.49 = 0.7 \times r\] so the common ratio is 0.7 and the 1st term is 1 then the you can use this information with n = 6 to find the sum

OpenStudy (anonymous):

I got .16807 as my answer, is that correct, @raster?